Hello ….… Two stars each of one solar mass ( = 2 x 10 ^30 kg) are approaching each other for a head on collision . When they are a distance of 10 ^9 km , their speeds are negligible . What is the speed with which they collide ? The radius of each other is 10 ^4 km . Assume the stars to remain undistorted until they colide .

Hello ….…
    Two stars each of one solar mass ( = 2 x 10 ^30 kg) are approaching each other for a head on collision . When they are a distance of 10 ^9 km , their speeds are negligible . What is the speed with which they collide ? The radius of each other is 10 ^4 km . Assume the stars to remain undistorted until they colide .

Grade:9

1 Answers

Vikas TU
7 years ago
Considering the total energy =>
0.5mv^2 + 0.5mv^2 – GMM/2R = -GMM/r
solving v^2 = GM(-1/r + 1/2R)
puttig the values given in qstn. we get,
v = root(6.66 x 10^12) => 2.58 x 10^6 m/s